### v2

# Few-Shot Recognition via Stage-wise Retrieval-Augmented Finetuning

Abstract
Few-shot recognition (FSR) aims to train a classification model with only a few labeled examples of each concept concerned by a downstream task, where data annotation cost can be prohibitively high. We develop methods to solve FSR by leveraging a pretrained Vision-Language Model (VLM). We particularly explore retrieval-augmented learning (RAL), which retrieves data from the VLM’s pretraining set to learn better models for serving downstream tasks. RAL has been widely studied in zero-shot recognition but remains under-explored in FSR. Although applying RAL to FSR may seem straightforward, we observe interesting and novel challenges and opportunities. First, somewhat surprisingly, finetuning a VLM on a large amount of retrieved data underperforms state-of-the-art zero-shot methods. This is due to the imbalanced distribution of retrieved data and its domain gaps with the few-shot examples in the downstream task. Second, more surprisingly, we find that simply finetuning a VLM solely on few-shot examples significantly outperforms previous FSR methods, and finetuning on the mix of retrieved and few-shot data yields even better results. Third, to mitigate the imbalanced distribution and domain gap issues, we propose Stage-Wise retrieval-Augmented fineTuning (SWAT), which involves end-to-end finetuning on mixed data in the first stage and retraining the classifier on the few-shot data in the second stage. Extensive experiments on nine popular benchmarks demonstrate that SWAT significantly outperforms previous methods by $>$6% accuracy.
1 Introduction
Few-shot recognition (FSR) aims to train a model with only a few examples per concept for downstream tasks, where data annotation can be costly. One motivational application is to train machines for automated data annotation by learning from a few examples of each concept provided by data annotation guidelines [37]. Motivated by this data annotation application, we focus on solving FSR, particularly by leveraging a foundational Vision-Language Model (VLM) and its pretraining data.
Status Quo. In the literature, FSR has served as a proxy task to study parameter-efficient finetuning (PEFT) [52, 58, 36, 66], and the robustness and generalization of pretrained models [27, 80, 81], etc. Consequently, most FSR methods focus on learning a few parameters, such as a classifier head [15, 76, 73, 35, 57, 60] or prompt tokens [80, 8] over a frozen pretrained backbone (e.g., VLM’s visual encoder). However, emphasizing learning fewer parameters without prioritizing recognition accuracy limits the performance of these methods in real-world applications such as automated data annotation [37]. Importantly, recent works [35, 57] show that the setup for a majority of FSR methods unreasonably assumes access to a much larger validation set for hyperparameter tuning, hindering their practical utility. Our work solves FSR with the realistic motivation of automated data annotation, adopting a rigorous setup without access to an unrealistic large val-set and optimizing for higher accuracy without limiting to learning only a few parameters.
Figure 1: A summary of few-shot recognition (FSR) benchmarking results over nine datasets. Somewhat surprisingly, although underexplored in the literature, finetuning the entire visual encoder solely on few-shot annotated data (green line) already outperforms previous methods [35, 57] by >3% accuracy! Yet, finetuning only on retrieved data by retrieval augmented learning (RAL, orange line) underperforms the state-of-the-art zero-shot methods [42]. This is due to that the retrieved data follows an imbalanced distribution and has domain gaps with the few-shot data (Fig. 3). By addressing these issues, our SWAT performs the best (red line), achieving $>$6% accuracy better than previous methods. Refer to Appendix Fig. 5 for detailed results on each of the nine datasets.
Motivation. Different motivations drive different FSR methods. Many existing FSR methods are factually motivated to study PEFT [52, 58, 36, 66] or learned models’ robustness and generalization [80, 81, 27]. This motivation has led to a “flawed” setup, as noted in [35, 57], where an unrealistically large validation set is used for hyperparameter tuning. In contrast, we solve FSR motivated by data annotation [37], avoiding large validation sets and prioritizing accuracy over learning a small number of parameters. This data annotation-driven perspective expands FSR research beyond the prevailing PEFT [15, 76, 35, 57, 60]. Interestingly, we find that finetuning a pretrained visual encoder on few-shot annotated examples already surpasses prior FSR methods without overfitting issues (Fig. 1)!
Furthermore, most FSR methods focus on adapting a pretrained Vision-Language Model (VLM) [35, 57, 60]. Differently, we also embrace VLM’s pretraining data, exploiting retrieval-augmented learning (RAL) [33, 36, 65, 22, 50, 42]. This strategy shines in zero-shot recognition, where state-of-the-art zero-shot methods adopt a VLM to retrieve relevant pretraining data for the concepts of interest and learn a classification model over such retrieved data [42]. Our work, for the first time, extend RAL to FSR. Despite its simplicity, we reveal interesting and novel challenges and opportunities. We elaborate on them below.
Challenges and Insights. Unlike prior FSR methods that learn a small number of parameters, we first explore full finetuning of the entire pretrained model (i.e., VLM’s visual encoder) on the few-shot examples. Somewhat surprisingly, this simple method significantly outperforms previous FSR methods (Fig. 1) with quite affordable computation costs (Table 5), owning to the small scale of few-shot data. In another line of research, state-of-the-art methods of zero-shot recognition (ZSR) adopt a strategy called retrieval-augmented learning (RAL), which retrieves relevant examples from VLM’s pretraining set and learns a classifier over them [42]. We test RAL for FSR by only using the retrieved data: both classifier learning (i.e., REAL-Linear [42]) and full finetuning barely surpass the non-RAL ZSR method REAL-Prompt [42] and underperform other FSR methods [57, 35], even trained with far more retrieved data. We identify two culprits (Fig. 3): (1) imbalanced distribution of retrieved data, and (2) domain gaps between the retrieved data and few-shot examples. To address the two issues, we propose our method, Stage-Wise retrieval-Augmented fineTuning (SWAT, cf. Fig. 2). SWAT first finetunes the VLM’s visual encoder on mixed retrieved and few-shot data, then re-trains the classifier solely on few-shot examples. This stage-wise training of our SWAT aligns with the decoupling learning strategy [26], which addresses the imbalanced distribution issue by first learning feature representations using all imbalanced data and then learning the classifier with balancing techniques. SWAT also mitigates the domain gap in a transfer learning manner, i.e., pretraining on larger source-domain data followed by finetuning on the target-domain data [34, 21, 19].
Contributions. We make three major contributions:
  1. 1.
    To solve few-shot recognition (FSR) for real-world applications, e.g., automated data annotation, we develop FSR methods focusing on better accuracy without restricting the number of learned parameters. This broadens FSR research space. For example, we find that finetuning a VLM’s visual encoder on few-shot examples already outperforms previous FSR methods by $>$3% (Fig. 1).
  2. 2.
    For the first time, we explore retrieval-augmented (RAL) learning for FSR, a technique well-studied in zero-shot recognition (ZSR). Despite its simple extension from ZSR to FSR, we identify novel and interesting challenges: the retrieved data follows an imbalanced distribution and has domain gaps with the few-shot examples.
  3. 3.
    We develop a simple method Stage-Wise retrieval-Augmented fineTuning (SWAT) that effectively mitigates the aforementioned imbalanced distribution issue and domain gaps, resulting in significantly better performance (>6%) than prior arts on nine benchmarks. Fig. 2 illustrates our SWAT and Fig. 1 summarizes our results.
Figure 2: Overview of our Stage-Wise retrieval-Augmented fine-Tuning (SWAT) for few-shot recognition (FSR). Consider the scenario where one wants to train a model on a few examples per concept concerned in data annotation guidelines. SWAT exploits a pretrained Vision-Language Model (VLM) and retrieves its pretraining data relevant to the concepts of interest. We observe that the retrieved data follows an imbalanced distribution and has domain gaps from the few-shot examples (Fig. 3). SWAT addresses the two issues jointly by first finetuning the VLM’s visual encoder on both retrieved and few-shot annotated data, then re-training the classifier only using the few-shot examples. Over nine FSR benchmarks, our SWAT achieves state-of-the-art performance, significantly outperforming previous methods by $>$6% accuracy (Fig. 1).
2 Related Work
Few-Shot Recognition (FSR). Early FSR methods leverage metric learning [58, 63, 3], meta learning [14, 48], transductive learning  [82, 25, 5] and graph neural networks (GNNs) [16, 51]. These approaches pretrain a model on a large meta-training set, then finetune it on task-specific few-shot examples, aiming for strong generalization. Recent works exploit Vision-Language Models (VLMs) pretrained on web-scale data [47, 23, 1] for FSR. With a frozen visual encoder, these methods harness the extraordinary zero-shot transfer capabilities of VLMs by learning better prompts [80, 79, 8, 27, 69, 72, 81] or lightweight adapters [15, 76, 73, 35, 57, 60, 77]. While utilizing pretrained VLMs achieve state-of-the-art performance on various benchmark datasets, these studies prioritize parameter-efficient finetuning (PEFT) with frozen backbones, even though finetuning more layers can yield better results [35, 57]. Furthermore, recent studies [35, 57] point out that these works unrealistically use large validation sets for hyperparameter tuning, limiting their real-world applicability. This focus on FSR as a proxy for PEFT [52, 58, 36, 66] or model robustness/generalizability [27, 80, 81] has shaped the current “flawed setup” and their limitations.
Contrasting to these methods, our work is driven by the real-world application of data annotation, prioritizing recognition accuracy over PEFT constraints. We leverage both VLMs and their pretraining data while strictly adhering to a validation-free setting. This realistic setup broadens FSR research. Notably, we find that simply finetuning a VLM’s entire visual encoder on few-shot examples already significantly outperforms previous FSR methods.
Vision-Language Models (VLMs) and Retrieval-Augmented Learning (RAL). VLMs, such as CLIP [47] and ALIGN [23], are pre-trained on web-scale data sampled from the open world. They learn to map image and text data into a common embedding space via contrastive learning on millions of image-text pairs. VLMs have demonstrated impressive zero-shot transfer capability across various downstream tasks. To further improve VLMs’ zero-shot accuracy, recent works [42, 36, 65, 22, 33] explore RAL, which retrieve external data (e.g., from the VLM’s pretraining dataset) relevant to the downstream task and then use such data to finetune the VLM or learn a classifier. Inspired by RAL’s success in achieving state-of-the-art zero-shot performance, for the first time, we extend RAL to few-shot recognition. We identify not only opportunities but also interesting challenges. We present a novel FSR method that addresses these challenges, achieving state-of-the-art few-shot recognition performance on standard benchmark datasets.
Data Issues in the Open World. Real open-world data poses various challenges to a machine-learned model. Two well-known challenges relevant to our work are imbalanced distribution of training data, and domain gaps between training and testing data. Imbalanced learning has been extensively studied through long-tailed learning [78, 26, 7, 71, 2]; recent work [42] points out that pretrained VLMs also suffer from imbalanced distribution, as pretraining data naturally follows a long-tailed distribution in the real open world. Indeed, we confirm that the retrieved data also follows an imbalanced distribution w.r.t concepts concerned in a downstream task (cf. Fig. 3). Among various methods that address the imbalanced or long-tailed distribution [78], a simple stage-wise training approach proves to be effective: first training a model using all the imbalanced data, and then retraining the classifier with balancing techniques [26]. Additionally, retrieved data has a clear domain gap with the few-shot annotated examples (Fig. 3), which impacts model performance when applied across domains. Transfer learning and finetuning on target-domain data [17, 41, 39] can effectively mitigate this gap. Our work shows that using VLM and RAL for FSR necessitates addressing these two issues jointly. Our proposed Stage-Wise retrieval-Augmented fineTuning (SWAT) effectively tackles both, significantly enhancing FSR performance.
Moreover, to tackle data scarcity and enhance model performance, data augmentation is widely adopted during model training [49, 56]. Some augmentation techniques are performed on individual data examples, e.g., random cropping and color jittering [9, 28, 64], while others mix multiple examples [6], e.g., MixUp [75] and CutMix [74]. Among various data augmentation techniques, CutMix proves to stabilize training, reduce overfitting, and improve model performance [74]. Later works extend CutMix by combining it with MixUp [43], smoothing patch boundaries [43], leveraging saliency masks [62, 29], and considering long-tailed distribution of training data [44]. Our approach augments data by mixing retrieved images with few-shot annotated ones, yielding substantial improvements in FSR.
Figure 3: Retrieved data shows domain gaps with downstream few-shot data and follows an imbalanced distribution. Left: we compare retrieved and few-shot annotated images for random categories from five benchmark datasets. The two sets of images exhibit clear domain gaps regarding image styles, background content, and even semantics, e.g., the animal with banded stripes in the DTD dataset. Right: retrieved data follows imbalanced distributions w.r.t concepts defined in different downstream tasks, as the VLM’s pretraining set does not contain sufficient examples for certain classes. Due to these two issues, leveraging the retrieved data to improve FSR presents significant challenges. Refer to Table 1 for quantitative justification of domain gaps, and Appendix Fig. 7 and 8 for additional examples of more datasets.
3 Our Proposed Methods
Our work focuses on solving few-shot recognition (FSR). We first outline the problem and evaluation setup, then introduce our methods, progressing from finetuning on few-shot data to our final approach, Stage-Wise retrieval-Augmented fineTuning (SWAT).
Problem Setup. Our FSR setup aims for high recognition accuracy on the concepts concerned by a downstream task. Our setup follows recent studies [77, 35, 57] that leverage Vision-Language Models (VLMs) for FSR. This setup also permits the use of VLM pretraining data [54, 53], available as “free lunch”, which has been exploited in zero-shot recognition [42, 36, 65, 22, 33] and is thus applicable to FSR. To support real-world applications like data annotation, this setup prioritizes recognition accuracy over limiting the number of parameter to learn, contrasting with many FSR works that treat FSR as a proxy for studying model robustness or generalization [27, 80, 81], and PEFT methods [52, 58, 36, 66].
Evaluation Setup. Our work aims to develop practical FSR solutions for real-world applications, such as learning from few-shot examples in annotation guidelines for automated data annotation. To support our study, we adopt a rigorous and realistic evaluation setup [57], which prohibits using a large validation set for hyperparameter tuning. This differs from most FSR works, which rely on artificially large validation sets for hyperparameter tuning, as noted in recent studies [57, 35]. Instead, we set hyperparameters to default values from the literature, including learning rate and weight decay, and follow  [57] to apply this same set of hyperparameters across all benchmark datasets.
Remarks. Previous FSR methods often evaluate the generalization of models in a base-to-novel setup, which splits an existing dataset into separate base and novel class sets. In this setup, models train on few-shot data from the “base classes” and are then evaluated on the “novel classes”. However, this base-to-novel setup does not align with the our motivation – the practical application of data annotation, where the annotation guidelines clearly define the target concepts of interest, without introducing extra novel classes. More importantly, as noted by [37], this base-to-novel setup appears to be artificial as VLMs have already seen examples of these “novel” classes in their Internet-scale pretraining dataset. Therefore, in this work, we follow [37] to evaluate FSR methods on the target classes concerned by a downstream task, without a base-novel split.
Table 1: We demonstrate the domain gaps between two data sources: retrieved data and few-shot annotated data from a downstream task. Following [61], we train a binary classification model on the two data sources to identify the data source of each image from a held-out testing set. We report the binary classification accuracy on the testing sets, which are created specifically for this study based on each of downstream dataset. Results validate the presence of domain gaps with $>$90% mean accuracy over different downstream-task datasets. Refer to Fig. 3 for a visual demonstration.
dataset Semi-Aves Flowers Aircraft EuroSAT DTD mean acc.
accuracy 86.8 90.0 94.4 98.5 81.4 90.2
3.1 Finetuning on Few-Shot Data
Recall that our FSR setup prioritizes high recognition accuracy for downstream tasks, different from prior works that focus on PEFT [35, 57]. This motivated us to first test finetuning the VLM’s visual encoder on the few-shot examples alone. To the best of our knowledge, this method has not been explored in the literature. Surprisingly, this embarrassingly simple method significantly outperforms previous FSR approaches by $>$3% (Fig. 1 and Table 2)! Although one may suspect overfitting when finetuning the whole visual encoder on the few-shot examples, our empirical results show that overfitting is not an issue (Table 2 and Fig. 4). This is likely because of the robust hyperparameters derived from other research lines [42, 35, 67, 32], including classifier initialization using text embedding (Table 12) and a smaller learning rate for updating the visual encoder (Appendix Section B). Moreover, owning to the small scale of few-shot data, learning more parameters (i.e., finetuning the entire pretrained visual encoder) is still quite efficient (Table 5).
3.2 Finetuning on Retrieved Data
Next, we explore retrieval-augmented learning (RAL) with the VLM and its pretraining data. RAL retrieves pretraining examples relevant to the concepts concerned by the downstream task, which has been extensively studied in zero-shot recognition [42, 36, 65]. However, to the best of our knowledge, RAL has not been explored for FSR. Below, we detail the approach and the interesting challenges of using RAL for FSR.
Retrieval Augmentation. RAL has proven effective in zero-shot recognition [36, 65, 42]. There exist various strategies for retrieving VLM’s pretraining data relevant to concepts of interest. Some methods match feature similarities between textual embeddings of target concepts and features of pretraining images or captions [36, 65], while others use string matching to find relevant text and retrieve corresponding images [42]. The latter approach is significantly more efficient, as it avoids the high storage costs of downloading all images and the computational expense of embedding calculations for each image. Importantly, the latter helps achieve better zero-shot recognition and is more efficient than the former [42]. Therefore, we adopt the string matching approach [42] to retrieve examples in this paper.
Novel Challenges. With the large amount of retrieved data, we finetune the visual encoder for classification. Surprisingly, finetuning on retrieved data underperforms the few-shot finetuning method, and performs even worse than the state-of-the-art zero-shot method (Fig. 1)! We identify two key culprits contributing to the inferior performance: (1) domain gaps between the retrieved data and the downstream-task few-shot examples, and (2) imbalanced distributions of retrieved data. We elaborate on the two issues below and present our solution in the next subsection.
Issue 1: Domain Gap. VLMs’ pretraining dataset consists of a large number of web-scraped image-text pairs from diverse sources. These images vary in styles, resolutions, and backgrounds, which are not aligned with the distribution of downstream-task images. We compare examples of retrieved images and few-shot annotated images for five image classification tasks in Fig. 3. We further quantitatively verify the domain gaps by following [61]. Concretely, we train a binary classification model to distinguish whether a given image originates from the downstream-task data source or the retrieved data source. The resulting classifier achieves $>$90% accuracy (Table 1), confirming the significant domain gaps between the two data sources.
Issue 2: Imbalanced Distribution. The retrieved data follows an imbalanced distribution, as shown in Fig. 3. Imbalanced distributions often negatively impact training. It is worth noting that VLM’s pretraining data naturally follows imbalanced or long-tailed distributions [42], because certain concepts are just rarer than others in the real world. Despite the extensive manual efforts to balance diverse concepts in recent work [70], the resulting pretraining dataset still exhibits long-tailed distributions.
Table 2: Comparison of SWAT with the state-of-the-art zero-shot and few-shot recognition methods. Our first method that finetunes the visual encoder solely on few-shot examples already outperforms prior arts significantly ($\sim$3%)! Following the rigorous setup without using a validation set for hyperparameter tuning [57], this few-shot finetuning method has no overfitting issues, as further justified in Fig. 4. Additionally, our final method, SWAT, achieves significant improvements by finetuning on a mix of retrieved and few-shot data. It outperforms the previously best few-shot recognition method CLAP [57] by $>$6% in accuracy. For reference, we copy results from [57] for the methods that use an artificially large validation set for hyperparameter tuning. Bold and underlined numbers mark the best and second best numeric metrics; superscripts denote improvements over CLAP [57]. Detailed results on each dataset are provided in Appendix Table 7.
strategy method venue & year mean accuracy of nine datasets
zero-shot methods prompting-based OpenCLIP [10] CVPR 2023 56.3
REAL-Prompt [42] CVPR 2024 62.6
retrieval-augmented REAL-Linear [42] CVPR 2024 64.8
4-shot 8-shot 16-shot
few-shot methods prompt-learning CoOp [80] IJCV 2022 61.0 64.6 68.4
PLOT [8] ICLR 2023 62.9 65.7 68.7
adapter-based CLIP-Adapter [15] IJCV 2023 59.6 64.5 68.1
TIP-Adapter [76] ECCV 2022 56.6 57.8 59.5
TIP-Adapter(f) [76] ECCV 2022 60.8 63.5 67.1
TaskRes(e) [73] ECCV 2022 63.5 67.1 69.9
CrossModal-LP [35] CVPR 2023 65.4 68.8 71.8
CLAP [57] CVPR 2024 66.9 70.0 72.9
finetuning-based few-shot finetuning ours 69.7+2.8 73.3+3.3 76.3+3.4
SWAT ours 73.5+6.6 76.0+6.0 78.2+5.3
3.3 Stage-Wise Retrieval-Augmented Finetuning
To address the issues of domain gaps and imbalanced distributions, we propose Stage-Wise retrieval-Augmented fineTuning (SWAT), a two-stage solution (Fig. 2). In the first stage, SWAT finetunes the VLM’s visual encoder end-to-end on a mixed set of retrieved and few-shot annotated data using cross-entropy loss. In the second stage, SWAT retrains the classifier exclusively on the few-shot data atop the finetuned encoder. Below, we explain why SWAT can jointly address these two issues.
Stage-wise learning addresses domain gaps. By training on a larger dataset (mixing the retrieved and few-shot annotated data), SWAT produces a more generalizable feature representation, finetuning upon which serves as a practice of transfer learning [34, 21, 19]. Through transfer learning, training on domain-shifted data (different from downstream-task data) can improve downstream task performance after finetuning [55, 68, 30]. This explains why SWAT effectively addresses domain gaps and significantly enhances performance on the downstream task.
Stage-wise learning addresses imbalanced distribution. Imbalanced distribution is a notorious challenge when training models in the real world, as extensively studied through the problem of long-tailed recognition [78, 26, 7, 71, 2]. In the literature, [24] shows a simple yet effective strategy that trains a model over imbalanced or long-tailed data in the first stage to obtain generalizable feature representations, and then retrains the classifier with balancing techniques in the second stage. Similarly, SWAT finetunes the VLM’s visual encoder on the imbalanced mix of retrieved and few-shot data to produce generalizable features. In the second stage, SWAT retrains the classifier on balanced few-shot examples, effectively mitigating the imbalanced distribution issue.
Data Augmentation. In our work, we adopt data augmentation techniques to enrich our few-shot annotated examples. Specifically, we apply the CutMix technique [74] which creates new images by cutting a random patch from one image and pasting it to another. CutMix has been shown to stabilize training and reduce overfitting [74]. In our experiments, CutMix significantly improves few-shot recognition accuracy with minimal computation overhead, while other more advanced augmentation techniques offer no additional gains but more computation overheads (Appendix Table 11).
Table 3: Comparison of the accuracy of common and rare classes with vs. without stage-2 classifier retraining. We define the rare classes as the 10% least frequent classes in retrieved data and the rest as the common classes. Results show that stage-2 classifier retraining clearly improves recognition accuracy on both common and rare classes across all methods, including finetuning on few-shot data only, on retrieved data only, and on mixed data with or without CutMix data augmentation. Importantly, the improvement for rare classes is more significant than for common classes, confirming that classifier retraining mitigates the issue of imbalanced distribution in the retrieved data. We report the mean accuracy over nine datasets using 16-shot examples. Accuracy improvements compared to the model after stage-1 finetuning are marked in superscripts and standard deviations across three runs with different random seeds are marked in subscripts. See detailed improvements for each dataset in Appendix Table 13.
data used in stage-1: finetuning mean accuracy of nine datasets
stage-1: finetuning stage-2: classifier retraining
common rare average common rare average
[…]
4 Experiments
We conduct extensive experiments to demonstrate that SWAT significantly outperforms previous few-shot recognition methods. We begin with the experiment setup, followed by benchmarking results. We also ablate important design choices with extensive analyses.
4.1 Experimental Setup
Datasets and Metrics. Motivated from the data annotation perspective, we study FSR through challenging real-world tasks where the data annotation requires domain expert knowledge, such as recognizing bird species, aircraft models, satellite images, etc. Specifically, we follow [13, 50] to use fine-grained recognition datasets where the VLM CLIP [47] struggles to recognize the nuanced attributes [50]. We select nine datasets, namely Semi-Aves [59], Flowers102 [40], FGVC-Aircraft [38], EuroSAT [38], DTD [11], OxfordPets [45], Food101 [4], StanfordCars [31], and ImageNet [12] (see detailed descriptions in Appendix Table 6). We evaluate SWAT using the OpenCLIP ViT-B/32 model [20] and retrieve images from its publicly available LAION-400M pretraining set [53, 54]. We report the accuracy averaged over nine datasets. The appendix includes results using other VLM architectures, on which our conclusions still hold.
Table 4: Ablation study on important components in our SWAT. Comparing to the state-of-the-art adapter-based few-shot method CLAP [57], our few-shot finetuning (FSFT) obtains $>$2% improvement! Naively adding retrieved data for finetuning brings small improvements $\sim$1due to the domain gaps and imbalanced distribution of the retrieved data. Applying CutMix provides an additional $\sim$2improvement, and retraining the classifier in the second stage brings a further $\sim$0.5increase. Note the improvements of classifier retraining is more significant for the rare classes ($>$2%), especially when the stage-1 finetuning data is imbalanced retrieved data ($>$11%), as shown in Table 3. Superscripts indicate the improvements of each component (relative to the corresponding row above). We report mean accuracies averaged over nine benchmark datasets. Detailed results are listed in Appendix Table 14.
method finetune model retrieve data apply CutMix retrain classifier mean acc. of nine datasets
4-shot 8-shot 16-shot
CLAP [57] 66.9 70.0 72.9
FSFT (ours) 69.4+2.5 72.7+2.7 75.1+2.2
70.8+1.4 73.0+0.3 75.3+0.2
73.0+2.2 75.2+2.2 77.3+2.0
SWAT (ours) 73.5+0.5 76.0+0.8 78.2+0.5
Compared Methods. We compare SWAT with state-of-the-art zero-shot and few-shot recognition methods. For zero-shot methods, we study OpenCLIP [20] and the recent methods: REAL-Prompt [42] and REAL-Linear [42]. For few-shot methods, we compare against prompt-learning-based methods CoOp [80] and PLOT [8], and adapter-based methods CLIP-Adapter [15], TIP-Adapter [76], TaskRes [73], Cross-modal Linear Probing [35] and CLAP [57]. Note that CLAP is the state-of-the-art FSR method that strictly adopts the realistic validation-free setup [57].
Implementations. For each dataset, we retrieve 500 images per class following [42]. We strictly follow the validation-free protocol as done in [57] and set the same hyperparameters in all datasets (see details in Appendix Section B). We conduct experiments using 4, 8, and 16 shots of data, randomly sampled from downstream training sets with three different seeds. We report their average accuracy. Note that in the motivational application, data annotation guidelines provide multiple visual examples. Therefore, we do not conduct experiments with fewer shots (e.g., 1 or 2-shot) in our experiments. We run all experiments on a Quadro RTX 6000 (24GB) GPU with 50GB storage for hosting retrieved data for all datasets. SWAT takes 2.5 hours for larger datasets like Semi-Aves (200 classes) and less than 4 minutes for smaller ones like EuroSAT (10 classes). Table 5 compares its compute cost with previous methods. We provide our code for open source in the supplementary material.
4.2 Benchmarking Results and Ablation Studies
SWAT significantly outperforms previous methods. Table 2 shows that simply finetuning on few-shot data readily outperforms the previous state-of-the-art methods without bells and whistles. Moreover, our SWAT surpasses previous methods by over 6% in accuracy. In challenging datasets like Semi-Aves, EuroSAT, Aircraft, our improvements are even more significant (10-25%, cf. Appendix Table 7). These results highlight the efficacy of our SWAT for FSR.
Stage-wise finetuning improves rare class accuracy. Table 3 shows overall improvements, with particularly significant gains on rare classes, are achieved through stage-wise finetuning and data augmentation. These results confirm the effectiveness of classifier retraining in handling imbalanced learning on retrieved data.
Ablation study. Table 4 shows that full finetuning leads to substantial accuracy gains ($>$2%). Simply adding retrieved data for finetuning brings slight gains (0.2-1.4%). Applying the CutMix data augmentation and classifier retraining as our SWAT brings more significant accuracy improvement ($>$2%). The results confirm that our SWAT effectively mitigates the two issues of domain gaps and imbalanced distributions of retrieved data.
Figure 4: Retraining the classifier on the few-shot data does not suffer from overfitting. We show the testing accuracies by retraining the classifier on 16 few-shot data at different epochs. We perform three runs of training with different random seeds. Results show that the testing accuracy does not decrease with more epochs and shows small standard deviations. We show the accuracy plots for other datasets in Appendix Fig. 11.
Longer training does not suffer from overfitting. Fig. 4 shows that testing accuracy does not decrease with more training epochs. This validates that learning on few-shot data with a strong pretrained backbone does not suffer from overfitting. This encourages future work in FSR to safely adopt the more rigorous validation-free setting, i.e., not using an artificially large validation set for hyperparameter tuning.
Further analyses. We conduct more analyses and provide them in Appendix. We summarize salient results here: properly filtering retrieved data improves performance further on some datasets (Table 9); CutMix [74] slightly outperforms other data augmentation methods with minimal computation overhead (Table 11); the string-match retrieval yields the best overall results, though different retrieval methods may excel on specific datasets (Table 10); initializing the classifier with text embedding before finetuning improves performance over random initialization (Table 12), partly explaining why finetuning on few-shot data does not suffer from overfitting; learning on more retrieved images shows diminishing performance gains (Table 15).
5 Discussions
Broader impacts. Our work explores practical few-shot recognition methods with a real-world motivation of data annotation, where annotation guidelines provide a few visual examples for each concept of interest [37]. Yet, similar to prior works, our methods exploit a pretrained VLM, using which may have negative impacts as VLMs could have learned biases from the Internet data. Moreover, our work exploits retrieval augmented learning to retrieve VLM’s pretraining data, which may also deliver biases. Lastly, our work has not evaluated the proposed method in the real-world data annotation application; we expect more challenges when applying our methods in the real world.
Table 5: Comparison of the compute cost of our SWAT with state-of-the-art few-shot recognition methods. We measure the GPU memory used and total training time using the Semi-Aves dataset (200 classes with 16-shot), and report mean accuracies averaged over nine datasets. Results show that simply finetuning on the few-shot data improves accuracy by 3% with slightly more training time and GPU memory cost than CLAP. In addition, SWAT improves accuracy significantly by >5% with very affordable retrieval and training costs. We highlight the accuracy of our methods in bold. Superscripts mark improvements compared to previous state-of-the-art CLAP [57].
method venue & yr mem. time mean acc.
CrossModal LP [35] CVPR’23 2 GB 2 mins 71.8
CLAP [57] CVPR’24 2 GB 2 mins 72.9
few-shot finetuning ours 5 GB 20 mins 76.3+3.4
SWAT retrieval ours 2 GB 1 hr 78.2+5.3
SWAT training ours 5 GB 2.5 hrs
Limitations and future work. We note several limitations and future directions. First, although our method achieves state-of-the-art performance by simply using hyperparameters reported in the literature, we believe that they are not optimal for FSR. Future work can explore human-in-the-loop intervention for hyperparameter tuning, e.g., manually constructing a validation set using retrieved data to tune hyperparameters. Second, as our method exploits concept names to retrieve pretrained data, future work can employ language models to generate richer descriptions or use the definitions provided in annotation guidelines for better retrieval. Third, despite our work shows that CutMix is effective in mitigating domain gaps and imbalanced distribution, it is worth exploring novel data augmentation methods to further address these issues. Lastly, although our method jointly solves the imbalance issue and domain gap of the retrieved data, future work can design experiments to decouple these two factors for further analysis and develop novel solutions.
6 Conclusions
We explore few-shot recognition (FSR) motivated from a data annotation perspective. This encourages one to solve FSR towards higher accuracy with a rigorous setup, unlike contemporary FSR works that focus on parameter-efficient finetuning a pretrained Vision-Language Model (VLM) in a flawed setup, which uses an unrealistically large validation set for hyperparameter tuning. We first examine an embarrassingly simple method that finetunes a VLM’s visual encoder on few-shot data. Surprisingly, it readily outperforms previous FSR methods. Then, we adopt the retrieval-augmented learning strategy, which retrieves VLM’s pretraining examples relevant to the concepts of interest to facilitate learning. We identify two interesting issues in the retrieved data: its imbalanced distributions, and its domain gaps with the few-shot examples. We propose a simple yet effective method, Stage-Wise retrieval-Augmented fineTuning (SWAT), to address the two issues. Across nine benchmark datasets, our SWAT outperforms previous methods by 6% accuracy on average, achieving the state-of-the-art FSR performance.
Supplementary Material
Outline
This document supports our main paper with detailed results and comprehensive analyses. The document is organized as below:
  • Section A. We provide a detailed summary of datasets used in our experiments.
  • Section B. We provide hyperparameters used in our work.
  • Section C. We report detailed results for each benchmark dataset.
  • Section D. We compare different retrieval and filtering methods.
  • Section E. We compare different data augmentation methods.
  • Section F. We compare different classifier initialization methods and provide detailed results of stage-wise training on each dataset.
  • Section G. We provide details of retrieved images for each dataset and analyze the impact of retrieval size.
  • Section H. We show retraining the classifier does not overfit for more datasets.
  • Section I. We provide our code and instructions.
A Summary of Datasets
We summarize the nine fine-grained datasets used in our experiments in Table 6. Following [35, 57], we sample few-shot data from the official training set and evaluate model performance on the official test set, except for ImageNet where we report performance on its validation set. We repeat each experiment for three times with three random seeds. Note that we strictly follow the validation-free protocol [57] that we do not use any validation data for hyperparameter tuning or model selection.
Table 6: Statistics of nine fine-grained datasets repurposed in our work. We list the number of images in the official training, validation, and test sets for each dataset. The protocol of few-shot recognition samples few-shot data from the official training set; we use them as our train set. We repeat the sampling and training for three times for each method with three random seeds. To evaluate methods, we repurpose their official test set as our test set (except on ImageNet where we use its official validation set as our test set). We benchmark methods on our test sets. Note again that we do not use any validation examples for model selection or hyperparameter tuning; instead, we strictly adhere to the realistic validation-free protocol for few-shot research [57].
dataset # cls official-train official-val official-test task
Semi-Aves [59] 200 3,959 2,000 4,000 recognize birds
Flowers [40] 102 4,093 1,633 2,463 recognize flowers
Aircraft [38] 100 3,334 3,333 3,333 recognize aircrafts
EuroSAT [18] 10 13,500 5,400 8,100 classify satellite images
DTD [11] 47 2,820 1,128 1,692 recognize textures
OxfordPets [45] 37 2,944 736 3,669 recognize pets
Food101 [4] 101 50,500 20,200 30,300 recognize food
StanfordCars [31] 196 6,509 1,635 8,041 recognize cars
ImageNet [12] 1,000 1.28M 50,000 N/A large scale recognition
B Hyperparameter Setting
We follow previous work in other lines [42, 35, 32] to set hyperparameters in our work. Specifically, for stage-1 end-to-end finetuning of SWAT, we follow suggestions from [32, 67] to use a smaller learning rate (1e-6) for updating the visual encoder and a larger learning rate (1e-4) for the linear classifier. We initialize the classifier weights using the text embedding following [42] (Table 12). For other hyperparameters, we adopt the values reported in [42, 35] which include the AdamW optimizer, a batch size of 32, weight decay of 1e-2, and a cosine-annealing learning rate schedule with 50 warm-up iterations. We do not do early stopping as we strictly follow the validation-free protocol [57]. Instead, we train for 50 epochs. The only exception is for ImageNet, we train for 10 epochs due to the large amount of retrieved data for its 1,000 classes. The temperature factor is learned during the finetuning process with an initial learning rate of 1e-4 and the same cosine-annealing learning rate schedule. For data augmentation, we mix retrieved data with few-shot data using CutMix [74], following [44] to sample the mixing ratio from a uniform distribution ($\alpha=1.0$ for beta distribution) and apply CutMix with a probability of 0.5. Our few-shot finetuning (FSFT) adopts the same set of training recipe.
For stage-2 classifier retraining, we use the same set of hyperparameters and follow the practice in [42] to train for 10 epochs with a fixed temperature 0.01. We initialize the classifier in stage 2 using the learned classifier weights from stage-1 end-to-end finetuning, following [32].
For other methods, we reimplement CrossModal Linear Probing [35] using the same hyperparameters as in stage-2 classifier retraining and train for 50 epochs. We obtain the results of CLAP [57] using OpenCLIP models with its default hyperparameters. For other baseline methods that originally uses an unrealistically large validation set for hyperparameter tuning, we copy their results from [57].
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 5: Comparison of SWAT with state-of-the-art zero-shot and few-shot methods. We show that simply finetuning the whole visual encoder on few-shot data (our few-shot finetuning, green line) outperforms previous FSR methods, while finetuning on retrieved data (orange line) underperforms zero-shot methods (e.g., ImageNet, EuroSAT, Food, DTD, and Stanford Cars) due to the large domain gap and imbalanced distributions of retrieved data. Our SWAT (red line) outperforms previous methods by $>$6% w.r.t accuracy over nine datasets, with significant improvements (20-30%) on challenging datasets like Semi-Aves and Aircraft. The results validates the effectiveness of our SWAT in mitigating the domain gap and imbalanced distribution issues. We provide detailed performance on each dataset in Table 7. For Flowers, EuroSAT, DTD, and Stanford Cars datasets, we show that SWAT can be further improved by 1-6% of accuracy with proper filtering on the retrieved data (red dashed line, Table 9).
Table 7: Detailed comparison of our SWAT and few-shot finetuning (FSFT) with state-of-the-art zero-shot and few-shot recognition methods using OpenCLIP ViT/B-32 model. SWAT significantly outperforms previous few-shot recognition methods by 6% across nine datasets. We highlight the best number in bold and underline the second best. Superscripts mark improvements compared to previous state-of-the-art FSR method CLAP [57].
shots strategy methods Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet average
0 prompting OpenCLIP [10] 8.4 68.2 17.1 51.1 53.5 88.7 77.2 79.2 63.0 56.3
REAL-Prompt [42] 43.4 76.0 18.0 56.9 59.2 88.7 77.1 80.6 63.6 62.6
retrieval-augmented REAL-Linear [42] 49.2 79.4 27.3 51.5 61.0 89.7 78.0 81.7 65.5 64.8
4 prompt-learning CoOp [80] 38.1 86.1 20.6 68.6 53.9 86.7 73.5 62.7 58.5 61.0
PLOT [8] 37.2 87.8 22.4 72.4 56.0 88.6 77.2 63.4 61.5 62.9
adapter-based CLIP-Adapter [15] 39.2 85.3 23.0 72.5 47.2 80.0 72.1 61.0 55.7 59.6
TIP-Adapter [76] 37.4 69.8 19.6 54.3 53.5 82.3 74.7 57.7 60.2 56.6
TIP-Adapter (f) [76] 42.4 74.4 21.9 66.8 58.0 85.5 75.3 61.1 61.5 60.8
TaskRes(e) [73] 43.2 89.4 25.9 73.0 58.4 84.6 74.5 64.7 58.0 63.5
CrossModal-LP [35] 29.1 88.9 25.1 74.8 62.2 88.3 76.7 80.7 63.2 65.4
CLAP [57] 34.0 90.1 28.0 74.7 63.0 87.0 76.7 84.9 64.0 66.9
finetuning-based FSFT (ours) 47.5+13.5 92.5+1.4 27.9-0.1 81.6+6.9 66.6+3.6 88.7+1.7 75.8-0.9 81.5-3.4 62.3-1.7 69.4+2.5
FSFT w/ CutMix (ours) 48.0+14.0 92.2+1.1 28.8+0.8 81.8+7.1 66.7+3.7 89.0+2.0 76.1-0.6 82.5-2.4 62.4-1.6 69.7+2.8
SWAT (ours) 58.5+24.5 90.6+0.5 55.7+27.7 83.2+8.5 58.3-4.7 91.3+4.3 77.3+0.6 81.1-3.8 65.8+1.8 73.5+6.6
8 prompt-learning CoOp [80] 42.0 91.3 26.6 77.1 59.7 85.4 71.6 67.6 60.4 64.6
PLOT [8] 41.4 92.4 26.2 78.2 61.7 87.4 75.3 67.0 61.9 65.7
adapter-based CLIP-Adapter [15] 41.2 91.9 27.9 78.5 61.4 83.4 72.1 66.8 57.0 64.5
TIP-Adapter [76] 39.8 73.8 19.4 62.3 51.5 82.3 73.9 57.6 59.4 57.8
TIP-Adapter (f) [76] 46.2 84.3 23.8 70.3 59.8 85.6 75.0 64.4 61.8 63.5
TaskRes(e) [73] 47.1 94.3 30.9 78.8 63.5 85.7 74.4 69.7 59.1 67.1
CrossModal-LP [35] 38.8 92.5 27.9 80.6 67.2 88.8 77.3 82.7 63.1 68.8
CLAP [57] 42.9 92.9 33.6 77.4 66.4 87.8 77.5 86.1 65.6 70.0
finetuning-based FSFT (ours) 51.2+8.3 95.4+2.5 33.1-0.5 90.3+12.9 71.0+4.6 89.3+1.5 76.0-1.5 83.5-2.6 64.4-1.2 72.7+2.7
FSFT w/ CutMix (ours) 52.3+9.4 95.2+2.3 35.4+1.8 89.4+12.0 70.6+4.2 89.6+1.8 77.0-0.5 85.3-0.8 64.8-0.8 73.3+3.3
SWAT (ours) 60.8+17.9 94.1+1.2 59.1+25.5 89.2+11.8 62.6-3.8 90.8+3.0 77.5+0.0 83.5-2.6 66.6+1.0 76.0+6.0
16 prompt-learning CoOp [80] 46.1 94.5 31.4 83.7 62.5 87.0 74.5 73.6 61.9 68.4
PLOT [8] 44.4 94.8 31.5 82.2 65.6 87.2 77.1 72.8 63.0 68.7
adapter-based CLIP-Adapter [15] 43.6 94.6 34.2 83.2 65.7 84.9 74.0 73.5 59.0 68.1
TIP-Adapter [76] 42.0 78.4 22.0 67.9 54.8 81.1 73.0 58.8 57.8 59.5
TIP-Adapter (f) [76] 50.1 91.2 29.3 76.6 64.6 85.4 74.7 69.6 62.3 67.1
TaskRes(e) [73] 48.5 96.1 36.5 83.7 65.9 86.3 75.4 75.4 60.9 69.9
CrossModal-LP [35] 46.8 95.5 32.4 85.2 71.9 89.1 77.5 84.7 63.1 71.8
CLAP [57] 49.2 94.8 39.1 81.7 69.9 88.4 78.5 87.8 67.1 72.9
finetuning-based FSFT (ours) 55.3+6.1 97.0+2.2 37.0-2.1 94.0+12.3 73.3+3.4 89.5+1.1 77.1-1.4 85.7-2.1 66.7-0.4 75.12.2
FSFT w/ CutMix (ours) 56.5+7.3 97.1+2.3 42.7+3.6 94.3+12.6 73.4+3.5 89.6+1.2 78.2-0.3 87.8+0.0 66.9-0.2 76.3+3.4
SWAT (ours) 63.1+13.9 96.4+1.6 62.4+23.3 92.6+10.9 66.3-3.6 91.6+3.2 78.3-0.2 85.4-2.4 67.6+0.5 78.2+5.3
Table 8: Detailed comparison of SWAT with state-of-the-art zero-shot and few-shot recognition methods using OpenCLIP ViT-B/16 model. Results show that SWAT achieves larger performance gains ($\sim$8%) than CLAP [57] with a larger backbone. We highlight the best number in bold and underline the second best. Superscripts mark improvements compared to previous state-of-the-art CLAP [57].
shots strategy methods Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet average
0 prompting OpenCLIP [10] 8.5 68.3 17.9 50.1 49.2 91.0 82.7 83.6 67.2 57.6
REAL-Prompt [42] 51.2 76.0 19.4 51.2 56.7 91.0 82.8 84.4 67.6 64.5
retrieval-augmented REAL-Linear [42] 57.1 80.3 29.2 46.8 60.3 91.4 83.3 85.5 69.8 67.1
4 adapter-based CrossModal-LP [35] 37.7 90.1 27.9 74.8 62.4 90.6 82.2 85.6 67.8 68.8
CLAP [57] 40.0 91.0 29.9 76.7 64.6 88.9 80.4 86.8 66.9 69.5
finetuning-based FSFT (ours) 57.7+17.7 93.6+2.6 33.0+3.1 85.5+8.8 69.1+4.5 91.4+2.5 81.9+1.5 86.1-0.7 67.4+0.5 74.0+4.5
FSFT w/ CutMix (ours) 58.8+18.8 93.4+2.4 33.4+3.5 83.4+7.7 68.6+4.2 91.8+2.9 82.7+2.3 87.0+0.2 67.8+0.9 74.1+4.6
SWAT (ours) 69.2+29.2 93.8+2.8 66.5+36.6 84.2+8.5 62.6-2.0 92.9+4.0 83.3+2.9 85.2-1.6 70.6+3.7 78.7+9.2
8 adapter-based CrossModal-LP [35] 49.4 93.6 32.5 81.8 67.8 90.9 82.9 87.4 68.0 72.7
CLAP [57] 49.1 93.4 36.1 79.0 67.7 89.6 81.5 88.4 68.5 72.6
finetuning-based FSFT (ours) 61.9+12.8 96.6+3.2 39.6+3.5 90.9+11.9 73.3+6.6 91.4+1.8 82.0+0.5 87.8-0.6 69.4+0.9 77.0+4.4
FSFT w/ CutMix (ours) 63.0+13.9 96.4+3.0 42.9+6.8 90.3+11.3 73.5+6.8 92.1+2.5 83.2+1.7 89.6+1.2 69.8+1.3 77.9+5.3
SWAT (ours) 71.4+22.3 96.5+3.1 69.1+33.0 88.8+4.6 66.3-1.4 93.2+3.6 83.8+0.5 87.2-1.2 71.5+3.0 80.9+8.3
16 adapter-based CrossModal-LP [35] 57.7 96.5 38.9 84.5 73.3 90.7 83.3 88.8 68.0 75.7
CLAP [57] 56.9 95.2 42.4 82.2 71.4 90.3 82.3 89.8 70.0 75.6
finetuning-based FSFT (ours) 66.3+9.4 98.0+2.8 45.6+3.2 94.1+11.9 75.8+4.4 91.5+1.2 82.5+0.2 89.7-0.1 70.2+0.2 79.3+3.7
FSFT w/ CutMix (ours) 67.3+10.4 98.2+3.0 51.2+8.8 94.2+12.0 76.1+4.7 92.3+2.0 84.0+1.7 91.3+1.5 72.1+2.1 80.7+5.1
SWAT (ours) 73.9+17.0 98.2+3.0 72.6+30.2 93.0+10.8 69.0-2.4 93.3+3.0 84.4+2.1 89.0-0.8 72.3+2.3 82.9+7.3
C Detailed Benchmarking Results
We compare our SWAT and few-shot finetuning (FSFT) with prior state-of-the-art zero-shot [20, 42] and few-shot recognition methods [35, 57] using the OpenCLIP ViT-B/32 model in Fig. 5 and list the detailed performance in Table 7. We also include the performance our few-shot finetuning without CutMix. Results show that SWAT outperforms previous FSR methods by $>$6% accuracy over nine datasets, with substantial gains (20-30%) on challenging datasets where prior FSR accuracy [35, 57] was below 50% (e.g., Semi-Aves and Aircraft). Additionally, SWAT with OpenCLIP ViT-B/16 model (Table 8) yields even higher gains of 8% over [57] across nine datasets.
Further analysis. Our experiments show that SWAT underperforms prior state-of-the-art FSR method [57] on DTD and Stanford Cars. We conjecture that this is due to the significant domain gaps in the retrieved data, finetuning on which could hurt model’s performance. This motivates us to apply a filtering technique to remove excessively out-of-domain retrieved data. Indeed, as shown in Table 9, applying proper filtering on the retrieved data significantly boosts the performance of SWAT, allowing it to outperform CLAP [57]. We also find filtering improves SWAT on other datasets, including Flowers and EuroSAT. We mark the improved SWAT’s performance on these datasets in Fig. 5.
Moreover, the improved SWAT still underperforms few-shot finetuning (FSFT) on DTD datasets. We hypothesize that the discrepancy is because of DTD’s strict data collection rules, which includes only images that are almost entirely filled with a texture [11]. In contrast, the retrieved images often have only part of region depicting the texture (Fig. 8 and 9). This suggests future work to explore better retrieval or filtering methods to find images that are better aligned with downstream distribution, e.g., by referring to the data collection rules provided in the data annotation guidelines. We explore different retrieval methods in Section D.
Table 9: Comparison of SWAT’s performance with prior state-of-the-art FSR method CLAP [57]. SWAT underperforms CLAP on DTD and Cars datasets due to the significant domain gaps. However, with proper filtering on retrieved images (by keeping the top-10 retrieved images for each class that are ranked by prompt-to-caption or T2T similarity and discarding others), SWAT outperforms CLAP. We show results of different retrieval size in Table 15. Subscripts mark the performance difference compared with CLAP [57].
dataset methods 4-shot 8-shot 16-shot
DTD CLAP [57] 63.0 66.4 69.9
SWAT 58.3-2.0 62.6-3.8 66.3-3.6
SWAT+filtering 63.5+0.5 69.1+2.7 72.9+3.0
Cars CLAP [57] 84.9 86.1 87.8
SWAT 81.1-3.8 83.5-2.6 85.4-2.4
SWAT+filtering 83.5-1.4 86.8+0.7 88.6+0.8
Table 10: Comparison of SWAT using different retrieval methods. We conduct experiments on six datasets by first conducting string matching following [42] to download images whose captions contain any of the concepts’ synonyms, then ranking the images using different text (few-shot concepts or database captions) and image (database images or few-shot images) features for selecting the images most relevant to downstream concepts. The top-ranking 500 images for each class are selected for running SWAT with 16 few-shot data. Results show that despite all methods outperforms random sampling by $<$1% in average accuracy, no single method is the best for all datasets. We highlight the best number in bold and underline the second best. We further explore adding a text-to-image filtering before text-to-text ranking to remove noisy images with image-to-FS-concept similarity less than 0.25. Results show that T2I filtering improves SWAT’s performance significantly, especially for DTD dataset (6% improvement). We show examples of T2I filtered images in Fig. 10.
retrieval/ranking method Semi-Aves Flowers Aircraft EuroSAT DTD Cars average
random sampling 62.8 96.0 62.2 92.6 64.9 84.7 77.2
text-to-text: FS-concept & DB-caption 63.4 96.4 62.7 93.0 65.8 85.4 77.8
image-to-image: FS-image & DB-image 63.0 97.1 62.8 92.7 64.9 84.9 77.6
image-to-text: FS-image & DB-caption 63.2 96.8 62.8 93.4 66.7 86.9 78.3
image-to-text: FS-concept & DB-image 62.9 96.8 63.3 93.4 65.7 83.7 77.6
text-to-image filtering (0.25) + text-to-text 63.8 97.5 62.6 93.7 71.6 85.8 79.2
D Analysis of Retrieval and Filtering Methods
Retrieval-augmented learning has been extensively studied for zero-shot recognition [36, 65, 42]. Previous work [36] utilizes text-to-text (T2T) or text-to-image (T2I) similarity to retrieve images relevant to each downstream concept. However, as noted by [42], such similarity-based retrieval requires significant storage for downloading all the source images (e.g., $>$10TB for LAION-400M) and high compute costs for computing image and text features (>250 T4 GPU hours). In addition, [65] points out the challenge of threshold selection in similarity-based retrieval: setting it too high includes irrelevant images, which can negatively impact training. Moreover, the proper threshold varies for different concepts, making it infeasible to search at scale. Given the above limitations, in this study, we adopt the string-matching-based retrieval by [42] in the following steps.
Step 1: we use string matching to retrieve images whose captions contain any of downstream concepts’ synonyms. This circumvents the large storage cost, as now we only need to download the text (60GB for LAION texts) for string matching and then the images with matching captions (50GB for all nine datasets). Additionally, [42] shows that using concept synonyms help retrieve diverse images which benefits the retrieval-augmented learning.
Step 2: we then rank the retrieved images based on prompt-to-caption (T2T) similarities and select the top-ranking 500 images for each downstream concepts. We compare different ranking methods using text and image features in Table 10. The results show that, despite all ranking methods outperform the random sampling, no single ranking method is the best across all datasets. This suggests future work to design retrieval methods customized to each downstream task.
In addition, we follow the practice in the curation of LAION dataset to apply text-to-image (T2I) filtering, excluding noisy retrieved images with T2I (few-shot concepts and retrieved images) similarities below 0.25. Results in Table 10 show that this filtering notably improves SWAT’s performance, especially on DTD dataset ($>$6%). We show examples of filtered noisy images in Fig. 10.
E Analysis of Data Augmentation Methods
We show examples of various mixed sample data augmentation (MSDA) methods in Fig. 6 and compare their performance using SWAT across five datasets (Semi-Aves, Flowers, Aircraft, EuroSAT and DTD) in Table 11. Results show that CutMix performs the best with minimal computation overhead, while SaliencyMix [62] performs similarly but incurs significant overhead due to the extraction of saliency maps.
Figure 6: Examples of different mixed sample data augmentation (MSDA) methods. We show two examples where the first row shows mixing the retrieved and few-shot images from the same class in Semi-Aves dataset [59], and the second row shows mixing images from different classes in the FGVC-Aircraft dataset [38]. These MSDA methods encourage the model to learn from small discriminative parts of the object or details in the background (e.g. part of a bird or airplane), thereby improving the performance. Compared to CutMix [74] and its variants (SaliencyMix [62], ResizeMix [46]), MixUp [75] augments data by simply interpolating two images, which may create unnatural artifacts that could confuse the model [74].
Table 11: Comparison of using different Mixed Sample Data Augmentation (MSDA) methods in SWAT. Compared with no mixing, all mixing methods increase accuracy by 1-2%. MixUp [75] slightly underperforms other CutMix variants, likely because it creates unnatural artifacts that could confuse the model [74]. We also find that randomly applying CutMix regardless of few-shot and retrieved images performs better than strictly cutting few-shot patches and pasting them into retrieved images (CutMix-strict), likely because doing so limits the diversity of data augmentation. By default, SWAT uses CutMix [74], which achieves the best performance and low computation overhead among all the compared MSDA methods. Bold and underlined numbers mark the best and second best numeric metrics; superscripts denote improvements over no mixing. See visual examples of different MSDA methods in Fig. 6.
MSDA method compute overhead mean accuracy of five datasets
4-shot 8-shot 16-shot
No mixing None 68.3 71.9 75.6
MixUp [75] Low 69.1+0.8 73.0+1.1 76.6+1.0
SaliencyMix [62] High 70.1+1.8 74.4+2.5 77.7+2.1
CMO [44] Med 69.9+1.6 74.1+2.2 77.1+1.5
ResizeMix [46] Med 69.6+1.3 74.1+2.2 77.2+1.6
CutMix-strict Med 70.1+1.8 73.8+1.9 77.6+2.0
CutMix [74] Low 70.5+2.2 74.2+2.3 77.8+2.2
Figure 7: Retrieved data follows imbalanced distribution for all nine datasets. The retrieved data for ImageNet, Food, DTD, and Pets datasets are less imbalanced than other datasets, likely because the concepts from these datasets naturally appear more frequently on the Internet [42].
F Further Analyses on SWAT
We compare different classifier initialization methods for SWAT (Table 12). Results show that initializing with text embedding yields better performance than random initialization. In addition. we show the detailed performance of classifier retraining for each dataset in Table 13. The rare classes of the Aircraft dataset show significant performance gains (>10%) after classifier retraining, demonstrating the efficacy of classifier retraining with few-shot data in mitigating domain gaps and imbalanced distribution. In addition, we include the detailed ablation of SWAT components on each dataset in Table 14. Results show that applying CutMix [74] and classifier retraining effectively mitigate the domain gap and imbalanced distribution problem, verifying the design of SWAT.
Table 12: Comparison of classifier initialization methods in SWAT. We compare the final test accuracy by initializing the classifier before stage-1 end-to-end finetuning in different ways. Initializing classifier weights with text embedding features leads to better performance than random initialization. [32] explains that using randomly initialized classifier weights to finetune the model can distort the features of pretrained model, leading to worse finetuning performance. Throughout this work, we use prompts in [42] to initialize classifier weights in SWAT.
classifier initialization mean accuracy of nine datasets
4-shot 8-shot 16-shot
random 72.7 75.1 77.5
text embedding [42] 73.6+0.9 76.1+1.0 78.2+0.7
Table 13: Detailed comparison of the accuracy of common and rare classes after stage-1 and stage-2 training. We define the rare classes as the 10% least frequent classes in retrieved data and the rest as the common classes. Results show that stage-2 classifier retraining clearly improves recognition accuracy on both common and rare classes in all methods, including finetuning on few-shot data only, on retrieved data only, and on mixed data with or without CutMix data augmentation. Importantly, the improvement on rare classes is more significant than that on common classes, confirming that classifier retraining mitigates the issue of imbalanced distribution in the retrieved data. We report the accuracy for each dataset using 16-shot examples.
data used in stage-1: finetuning stage classes Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet average
few-shot only (balanced) stage-1 finetuning common 56.1 97.6 43.2 94.9 73.5 90.6 78.8 88.6 68.0 76.8
rare 63.5 100.0 34.6 87.1 76.7 84.9 74.8 84.4 56.8 73.6
average 56.9 97.4 42.4 94.1 73.9 90.0 78.4 88.0 66.9 76.4
stage-2 classifier retraining common 56.0 97.5 47.3 95.0 72.9 90.0 78.9 88.5 67.1 77.0
rare 63.8 100.0 46.4 87.2 76.7 86.9 74.5 83.3 57.3 75.1
average 56.8 97.4 47.2 94.3 73.3 89.7 78.4 87.9 66.1 76.8
retrieved only (imbalanced) stage-1 finetuning common 56.2 84.4 52.5 30.4 52.4 90.8 76.0 78.1 62.5 64.8
rare 15.0 54.4 10.2 0.0 61.1 85.5 73.3 51.8 46.4 44.2
average 52.1 81.6 48.3 27.9 53.3 90.3 75.7 75.3 60.9 62.8
stage-2 classifier retraining common 60.0 90.2 57.5 32.2 54.6 90.9 76.8 82.9 64.7 67.8
rare 36.9 77.8 33.7 0.0 62.8 86.6 74.2 66.8 58.8 55.3
average 57.7 88.6 55.1 29.4 55.4 90.5 76.6 81.2 64.1 66.5
retrieved + few-shot stage-1 finetuning common 61.4 94.6 57.4 93.4 62.3 91.4 77.9 81.8 64.8 76.1
rare 49.4 96.8 26.2 87.5 69.4 87.4 75.9 68.8 52.7 68.2
average 60.2 94.7 54.3 92.8 63.1 91.0 77.7 80.3 63.6 75.3
stage-2 classifier retraining common 61.6 95.4 60.6 93.4 62.8 91.3 78.0 84.0 65.7 77.0
rare 52.2 98.0 44.3 87.5 68.9 87.1 76.0 73.1 57.6 71.6
average 60.6 95.4 59.0 92.8 63.5 91.0 77.8 82.8 64.9 76.4
retrieved + few-shot w/ CutMix stage-1 finetuning common 63.7 96.4 61.3 93.4 64.8 91.5 78.3 83.9 68.3 78.0
rare 55.8 100.0 34.7 83.9 72.2 89.2 77.4 78.0 56.1 71.9
average 62.9 96.3 58.7 92.5 65.6 91.3 78.2 83.2 67.1 77.3
stage-2 classifier retraining common 64.0 96.4 63.7 93.7 65.6 91.9 78.4 86.1 68.3 78.7
rare 54.9 100.0 50.9 82.0 72.2 88.6 77.5 79.9 61.2 74.1
average 63.1 96.4 62.4 92.6 66.3 91.6 78.3 85.4 67.6 78.2
Table 14: Ablation study on important components in our SWAT. We show the detailed performance improvements by each component for each dataset in our SWAT. Finetuning on simply combined retrieved and few-shot data underperfroms finetuning solely on few-shot data (8-shot and 16-shot, with or without CutMix), due to the large domain gap and imbalanced distribution in retrieved data. However, further applying CutMix and classifier retraining improves the test accuracy significantly, confirming their effectiveness in mitigating the domain gap and imbalanced distributions. We also compare the performance of few-shot finetuning with and without CutMix data augmentation. The results indicate more few-shot data yields more improvements, likely due to stronger data augmentation.
shots method finetune model retrieve data apply CutMix retrain classifier Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet average
4 CLAP [57] 34.0 90.1 28.0 74.7 63.0 87.0 76.7 84.9 64.0 66.9
FTFS (ours) 47.5 92.5 27.9 81.6 66.6 88.7 75.8 81.5 62.3 69.4
FTFS (ours) 48.0 92.2 28.8 81.8 66.7 89.0 76.1 82.5 62.4 69.7
54.7 89.7 50.1 80.2 56.3 90.7 76.4 76.9 61.8 70.8
57.9 90.2 53.8 83.2 58.7 91.0 77.2 79.8 65.2 73.0
SWAT (ours) 58.5 90.6 55.7 83.2 58.3 91.3 77.3 81.1 65.8 73.5
8 CLAP [57] 42.9 92.9 33.6 77.4 66.4 87.8 77.5 86.1 65.6 70.0
FTFS (ours) 51.2 95.4 33.1 90.3 71.0 89.3 76.0 83.5 64.4 72.7
FTFS (ours) 52.3 95.2 35.4 89.4 70.6 89.6 77.0 85.3 64.8 73.3
57.3 91.9 52.4 87.0 59.2 91.1 76.8 78.9 62.5 73.0
60.6 93.7 55.7 89.1 61.8 90.8 77.6 81.3 65.8 75.2
SWAT (ours) 60.8 94.1 59.1 89.2 62.6 90.8 77.5 83.5 66.6 76.0
16 CLAP [57] 49.2 94.8 39.1 81.7 69.9 88.4 78.5 87.8 67.1 72.9
FTFS (ours) 55.3 97.0 37.0 94.0 73.3 89.5 77.1 85.7 66.7 75.1
FTFS (ours) 56.5 97.1 42.7 94.3 73.4 89.6 78.2 87.8 66.9 76.3
60.2 94.7 54.3 92.8 63.1 91.0 77.7 80.3 63.6 75.3
62.9 96.3 58.7 92.5 65.6 91.3 78.2 83.2 67.1 77.3
SWAT (ours) 63.1 96.4 62.4 92.6 66.3 91.6 78.3 85.4 67.6 78.2
Table 15: Impact of retrieval size (number of images per class) on the performance of SWAT. We show the performance of SWAT on each dataset using different numbers of retrieved images. We highlight the best number in bold and underline the second best. Importantly, we find that retrieving 10 images per class works the best for Flowers, EuroSAT, DTD, and Cars datasets. This is probably because LAION-400M contains limited images that matches these downstream concepts and simply retrieving more will include more noisy images and more imbalanced distributions, which hurt the training performance. We list the performance of previous state-of-the-art few-shot recognition method CLAP [57] in the table for comparison with our SWAT.
shots Retrieval size Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet average
4 CLAP [57] 34.0 90.1 28.0 74.7 63.0 87.0 76.7 84.9 64.0 66.9
10 52.4 91.8 37.0 84.7 63.5 89.3 75.9 83.5 64.8 71.4
100 57.4 90.7 47.0 82.1 62.1 89.9 76.6 83.5 66.1 72.8
300 58.7 91.4 54.1 82.2 59.3 91.1 77.1 81.7 65.9 73.5
500 58.5 90.6 55.7 83.4 58.3 91.3 77.3 81.1 65.8 73.6
1,000 58.3 89.6 58.1 84.1 57.7 91.4 76.2 81.1 65.2 73.5
8 CLAP [57] 42.9 92.9 33.6 77.4 66.4 87.8 77.5 86.1 65.6 70.0
10 55.7 95.2 42.2 90.0 69.1 89.4 76.9 86.8 65.8 74.6
100 59.2 94.6 49.9 88.6 65.2 90.2 77.2 85.3 67.0 75.2
300 60.6 94.3 56.5 89.3 63.1 90.9 77.6 83.9 67.3 75.9
500 61.3 94.1 59.1 88.7 62.6 91.5 77.6 83.5 66.6 76.1
1,000 60.9 92.9 60.6 88.9 59.8 91.4 76.7 83.6 66.2 75.7
16 CLAP [57] 49.2 94.8 39.1 81.7 69.9 88.4 78.5 87.8 67.1 72.9
10 58.4 97.0 48.6 94.0 72.9 89.6 78.5 88.6 66.9 77.2
100 61.8 96.8 54.5 93.4 69.4 90.2 78.6 87.1 67.9 77.7
300 63.2 96.8 60.8 93.1 67.0 91.3 78.6 86.0 67.8 78.3
500 63.1 96.4 62.4 92.9 66.3 91.6 78.3 85.4 67.6 78.2
1,000 63.6 96.4 64.2 93.0 63.0 91.8 77.2 85.6 67.2 78.0
G Analysis of Retrieved Data
We show the imbalanced distribution of retrieved data for all nine datasets in Fig. 7. We report the total number of retrieved images per dataset with increasing retrieval size (images per class) in Table 16. With increasing retrieval size, the total number of retrieved images increases less significantly due to the limited presentation of many downstream concepts in the pretraining datasets (e.g. LAION [53, 54]). To address this issue, we suggest future work to retrieve relevant images from diverse data sources, e.g. other datasets or the Internet [33].
Additionally, we compare SWAT’s performance on different retrieval size in Table 15. Results show that SWAT saturates at 500 images per class for 4-shot and 8-shot cases, and at 300 for 16-shot. Notably, for Flowers, EuroSAT, DTD, and Cars, retrieving 10 images per class yields the best results, likely due to improved data balance and the removal of noisy images (Fig. 10). Future work can study how to enhance the balance and quality of retrieved data. Fig. 8 shows more examples of retrieved images for each dataset.
Figure 8: Comparison of downstream few-shot data with retrieved pretraining images (from LAION-400M [53]) for nine fine-grained datasets. We present more examples of retrieved images for randomly selected classes from each dataset. Compared to downstream few-shot images, the retrieved data exhibits diverse styles, backgrounds, resolutions, and even semantics, demonstrating significant domain gaps.
Figure 9: Visual comparison between downstream DTD images and the retrieved images (from LAION-400M [53]) for various DTD concepts. Clearly, large domain gap exists between the two data resources regarding styles, backgrounds, semantics, etc. In addition, the retrieved images only have a partial region depicting the texture, contrasting to the few-shot image which are “almost entirely filled with a texture” according to DTD’s strict data collection rules [11]. We suggest future work to explore better retrieval methods that are closely aligned with downstream data distribution, e.g., by referring to the data collection/annotation rules provided in the data annotation guidelines of a downstream task.
Figure 10: Examples of noisy retrieved images (from LAION-400M [53]) filtered by T2I thresholding. We show that string-matching-based retrieval (by searching image captions that contain any concept synonyms) can retrieve noisy images that could compromise the learning of downstream concepts, e.g., the bird eggs or the distribution map of bird species (first row). Using text-to-image (T2I) filtering helps remove such noisy images, and improve the performance of SWAT (Table 10). We choose a T2I threshold of 0.25 for our experiment, similar to that used in the curation of LAION [53, 54]. We highlight the T2I cosine similarity under each image and concept synonyms in the each image caption.
Table 16: Total number of retrieved images for each dataset under different retrieval size. With a larger retrieval size (number of retrieved images per class), we observe a diminished increase in the total number of retrieved images. This is because many downstream concepts have limited presence in the pretraining set (LAION-400M [53, 54]). See the imbalanced distribution of each dataset in Fig. 7.
imgs per class Semi-Aves Flowers Aircraft EuroSAT DTD Pets Food Cars ImageNet
10 1,940 1,002 1,000 71 470 370 1,010 1,939 9,989
100 15,687 9,376 9,120 530 4,700 3,700 10,100 18,494 98,753
300 34,685 25,140 21,774 1,330 14,100 11,100 30,241 51,251 288,532
500 47,006 39,465 30,429 1,871 23,364 18,460 49,914 80,648 471,876
1,000 67,418 71,332 44,519 2,387 45,978 36,105 96,697 147,568 901,902
H Analysis of Classifier Retraining
In Fig. 11, we show the final test accuracy after retraining the classifier across varying epoch numbers. For all datasets, accuracy remains stable with additional epochs. The small standard deviations across three runs with different random seeds confirm that stage-2 classifier retraining with few-shot data does not suffer from overfitting.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 11: Retraining the classifier on the few-shot data does not suffer from overfitting. We show the final test accuracies by retraining the classifier on the few-shot data for different epoch numbers. For each dataset, we perform three runs of training with different random seeds. Results show that testing accuracy remains stable with more epochs and shows small standard deviations, indicating classifier retraining does not suffer from overfitting.
I Code and Instructions
We release open-source Python code at https://github.com/tian1327/SWAT.
Requirements. Running our code requires some common packages. We installed Python and most packages through Anaconda. A few other packages might not be installed automatically, such as clip, open_clip_torch, img2dataset, torchvision, and PyTorch, which are required to run our code. We provide detailed instructions for building environment in file ENV.md. Below are the versions of Python and PyTorch used in our work.
  • Python version: 3.8.19
  • PyTorch version: 2.0.1
We suggest assigning $>$50GB storage space and $>$5GB GPU RAM to reproduce our experiments.
License. We release open-source code under the MIT License to foster future research in this field.
Instructions. We provided detailed step-by-step instructions for running our code in the following markdown files.
  • ENV.md
    Create conda environment and install required packages.
  • DATASETS.md
    We provide detailed steps for setting up the benchmarking datasets and sampling few-shot data from the official training sets with three random seeds.
  • RETRIEVAL.md
    We provide step-by-step instructions on how to use string-matching [42] to retrieve relevant images from OpenCLIP’s pretraining dataset LAION-400M [53, 54]. Examples of different ranking and filtering methods for selecting the images that are most relevant to downstream concepts are also provided.
  • README.md
    We provide instructions on how to run the provided code for few-shot finetuning (FSFT) and SWAT. In addition, we provide guidelines on how to reproduce the baseline methods Cross-Modal Linear Probing [35] and CLAP [57].